OpenStudy (anonymous):

log base 5^(x+4)+log base 5^(x+8)=1

OpenStudy (anonymous):

log base 5 ^(x+4)(x + 8)=1 in exponential form 5^1=(x+4)(x+8)

OpenStudy (anonymous):

the answers should be -3 but how do i get there?